OpenStudy (anonymous):

solve the equation: -6m+7=-7m+10

OpenStudy (benlindquist):

add your like terms first

OpenStudy (anonymous):

so add 6m to the -7m right?

OpenStudy (kendricklamar2014):

Step 1. Add 7m to both sides: \[−6m+7=−7m+10\] \[m+7=10\] Step 2. Subtract 7 from both sides:\[m+7−7=10−7\] \[m = 10 - 7 +?\] What is 10 - 7? @dazzling_nightmares

OpenStudy (anonymous):

3 the answer is 3

OpenStudy (kendricklamar2014):

Yes, the answer is 3!

OpenStudy (anonymous):

thank you

OpenStudy (kendricklamar2014):

Welcome :)